OpenStudy (anonymous):

Find the mean and standard deviation of each data set. Round to the nearst hundredth. 1. 4 8 5 12 3 9 5 2 2. 102 98 103 86 101 110 3. 8.2 11.6 8.7 10.6 9.4 10.1 9.3

OpenStudy (amistre64):

do you have to do them by hand? or can you use a calculator? or excel even?

OpenStudy (anonymous):

@amistre64 We can use a calculator but we have to show the work. I'd just like to know the problems step-by-step :)

OpenStudy (amistre64):

the mean of a set is the sum of the parts, divided by how many parts there are.

OpenStudy (anonymous):

So the mean is 6, for the first one.

OpenStudy (amistre64):

the standard deviation of a population set .... is the square root of variance variance is a little tricky to put into words, but is basically the mean of the squares of the "zeroed out" set

OpenStudy (amistre64):

sum{4 8 5 12 3 9 5 2} = 48 there are 8 data points ... mean is 48/8 = 6 yes

OpenStudy (amistre64):

now, the variance is simple enough to perform, just zero out set: subtract the mean from all the elements

OpenStudy (amistre64):

{4 8 5 12 3 9 5 2} - 6 -2 2 -1 6 -3 3 -1 -4 square these results and take the mean of the squares 4 4 1 36 9 9 1 16 sum{4 4 1 36 9 9 1 16} = 80 variance = 80/8, std dev = sqrt(variance)

OpenStudy (amistre64):

to "show your work" you prolly have to structure it into a table ... and prettify it

OpenStudy (anonymous):

So is the final answer 9?

OpenStudy (amistre64):

the final answer ... mean = 6 sd = sqrt(8)

OpenStudy (anonymous):

@amistre64 So is the answer "sqrt(8) = 3" or is it just "sqrt(8)"

OpenStudy (anonymous):

I got 3 for standard deviation and 6 for mean, can someone confirm this for me?